OpenStudy (anonymous):

Find the equation of the tangent lines to the parabola y=x^2+x that pass through the point (2, -3) . Sketch the curve and the tangents.

OpenStudy (freckles):

tangent line to f(x)=x^2+x at x=a is given by \[y-f(a)=f'(a)(x-a)\]

OpenStudy (freckles):

So you will need to find f'(x) first

OpenStudy (freckles):

can you do that?

OpenStudy (anonymous):

is it y'=2x+1??

OpenStudy (freckles):

yeah now what is f(a) and f'(a)

OpenStudy (freckles):

\[f(x)=x^2+x \\ \text{ so } \\ f(a)=? \\ \text{ and } f'(x)=2x+1 \\ \text{ so } f'(a)=?\]

OpenStudy (anonymous):

f(a)= a^2+a f'(a)=2a+1

OpenStudy (freckles):

great so we have this so far: \[y-f(a)=f'(a)(x-a) \\ y-(a^2+a)=(2a+1)(x-a)\]

OpenStudy (freckles):

now we know a point on this line

OpenStudy (freckles):

it was given to us

OpenStudy (freckles):

we wanted the point (2,-3) on this line so we can find a such that will happen by replacing x with 3 and y with -3 and solving for a

OpenStudy (freckles):

\[y-f(a)=f'(a)(x-a) \\ y-(a^2+a)=(2a+1)(x-a) \\ -3-(a^2+a)=(2a+1)(3-a)\] see if you solve this for a it should just be a quadratic

OpenStudy (anonymous):

i have a question, should it be x=2, not x=3

OpenStudy (freckles):

yes I'm dyslexic today

OpenStudy (freckles):

actually was one of them negative

OpenStudy (freckles):

x=2 and y=-3

OpenStudy (anonymous):

it will be -(a-5)(a-1) then a=1,5

OpenStudy (freckles):

hmm...

OpenStudy (freckles):

I got similar values one is just a bit different in sign then one you got

OpenStudy (anonymous):

i mean -(a-5)(a+1) a=5, -1

OpenStudy (anonymous):

oops

OpenStudy (freckles):

great work

OpenStudy (freckles):

that is exactly what i got

OpenStudy (freckles):

now we have to a few more things

OpenStudy (freckles):

it asked for equation of tangent line for which it was suppose to go through (2,-3)

OpenStudy (freckles):

\[y-f(a)=f'(a)(x-a) \\ y-(a^2+a)=(2a+1)(x-a) \\ \text{ replace } a \text{ with } 5 \text{ for one possible equation } \\ \text{ then replace } a \text{ with -1 } \text{ for another possible equation }\]

OpenStudy (freckles):

Are you cool with that?

OpenStudy (freckles):

for example I will give you equation: \[a=5 \\ y-(5^2+5)=(2 \cdot 5+1)(x-5) \\ y-(25+5)=(10+1)(x-5) \\ y-30=11(x-5)\] this equation for the tangent line at x=5 is in point-slope form

OpenStudy (anonymous):

still working on it

OpenStudy (freckles):

now take the same equation I just used and replace a's with -1

OpenStudy (freckles):

I don't know what the required form is like slope-intercept or point slope form

OpenStudy (anonymous):

i got y=x+1

OpenStudy (freckles):

\[a=-1 \\ y-(a^2+a)=(2a+1)(x-a) \\ y-([-1]^2+[-1])=(2 \cdot [-1]+1)(x-[-1]) \\ y-(1-1)=(-2+1)(x+1) \\ y-0=(-1)(x+1) \\ y-0=-1(x+1) \\ y=-(x+1)\] check it one more time just and case I made a mistake

OpenStudy (anonymous):

it looks right

OpenStudy (freckles):

your equation or mine

OpenStudy (anonymous):

yours, i forgot a negative in mine

OpenStudy (freckles):

ok just checking to see if you agree with my steps and also I'm not wearing my glasses lol

OpenStudy (freckles):

do you have any questions on this problem?

OpenStudy (freckles):

you have two answers

OpenStudy (anonymous):

how do you show it as "11x-y-25=0" and "x+y+1=0" (this is the answer from my textbook)?

OpenStudy (anonymous):

nevermind, i got it. @freckles Thank you for helping me! :D

OpenStudy (freckles):

alrighty :)

OpenStudy (freckles):

your mention didn't show through for some reason